OpenStudy (azureilai):

Can somone explain when the horizontal asymptotes can be crossed? I looked online, and the explanations are just confusing. Please help.

OpenStudy (anonymous):

Difficult question to answer here. Involves the calculus.

OpenStudy (anonymous):

yes, they can be crossed i can give two easy examples

OpenStudy (anonymous):

The question is WHEN can they be crossed. That is a complicated question to answer here.

OpenStudy (anonymous):

\[\frac{1}{x^2-1}\] as a horizontal asymptote at \(y=0\) but it crosses the \(x\) axis

OpenStudy (azureilai):

Well, I am working on a graph involving asymptotes right now, and the horizontal asymptote is apparently crossed (When I graphed it it showed) However, I am confused on how to algebraically see whether the horizontal asymptote is crossed or not.

OpenStudy (azureilai):

@satellite73 Can you explain why please?

OpenStudy (anonymous):

ok i lied, my example was wrong here is a better one \[y=\frac{x}{x^2-1}\]

OpenStudy (anonymous):

that has a horizontal asymptote at \(y=0\) aka the \(x\) axis

OpenStudy (anonymous):

but it clearly crosses the \(x\) axis because if \(x=0\) you get \(y=0\)

OpenStudy (anonymous):

horizontal asymptote is what the function does as \(x\to \infty\) or \(x\to -\infty\) what it does in the middle is not relevant so it can cross it

OpenStudy (azureilai):

Oh ok, so as long as it is in the middle, it can cross the Horizontal asymptote (under specific circumstances)?

OpenStudy (anonymous):

yes it can cross it in fact, although you have probably not seen these, it can cross infinitely often

OpenStudy (azureilai):

Ok thank you. That cleared up some misunderstandings I had.

OpenStudy (anonymous):

here is an example \[y=\frac{\sin(x)}{x}\] http://www.wolframalpha.com/input/?i=sin%28x%29%2Fx

OpenStudy (anonymous):

the vertical asymptote cannot be crossed